# Statistical Control of Relaxation and Synchronization in Open Anyonic Systems

### v2

@@ I Introduction
We recently investigated noise-induced synchronization in open quantum systems consisting of two coupled quantum oscillators interacting with a common, correlated dissipative environment [14, 15, 16]. In the limit of perfect environmental correlation ($\xi=\pm 1$), the oscillators exhibit long-lived phase synchronization, particularly when they are spectrally near-resonant. The presence of nonzero quantum discord between subsystems indicated that correlated environmental noise can generate persistent quantum coherences even in the absence of direct coupling. We further showed that two spins can synchronize their relative phases through shared stochastic driving, akin to the escapement mechanism that regulates motion in classical pendulum clocks.
@@ II Theoretical Development
We begin by developing a Lindblad framework for fractional-statistics oscillators, where the algebra of creation and annihilation operators is deformed by an exchange phase $\theta$ that interpolates continuously between bosonic and fermionic limits. For a single anyon oscillator, the operators $a^{\dagger}$ and $a$ satisfy the $q$-deformed commutation relation:
@@ II.1 Anyon master equation
We first consider the dynamics of a single anyon oscillator coupled to a dissipative bath using the adjoint Lindblad formalism. The system is governed by the Hamiltonian $H=\omega a^{\dagger}a$, where $a$ and $a^{\dagger}$ are the anyon annihilation and creation operators, respectively. Dissipative coupling is introduced via two Lindblad jump operators: $L_{\text{em}}=\sqrt{\gamma(n_{\theta}+1)}\,a$ for spontaneous and stimulated emission, and $L_{\text{abs}}=\sqrt{\gamma n_{\theta}}\,a^{\dagger}$ for absorption. The thermal occupation $n_{\theta}$ may be complex due to the anyonic exchange phase $\theta$.
In the Supporting Information we show how to derive the Lindblad equations of motion for anyonic operators $a$ and $a^{\dagger}$, which describe the coherences within the anyonic oscillator. Taking the expectation value over a thermal ensemble and including both thermal noise and statistical interference, the full coherence equation becomes:
@@
which remains finite at $\theta=0$ (c.f. Fig. 1). For bosonic systems ($\theta=0$), the statistical contribution vanishes, leaving only the thermal component. As the anyon phase increases towards the fermonic regime, the stastistical contribution increase monotonically until its maximum value at $\theta=\pi$ for a purely fermonic system. Consequently, moving from a purely fermionic system towards a fractionally quantized system results in a decrease in relaxation rate of the oscillator at a given temperature.
@@ II.2 Anyon double oscillator
with normal mode frequencies $\omega_{\pm}=\omega\pm J\cos(\theta)$. Following along the lines of our previous works[14, 15, 16] , (c.f. Appendix B) we define Lindblad jump operators in terms of the deformed normal mode operators as
@@
with explicit coefficients $\lambda_{k}^{(\pm)}$ for the emission and absorption channels given by Eqs.46-53 in the appendix. Using these, we derive the equations of motion for the deformed anyonic operators which we write in matrix form as
@@
where the effective evolution matrix is given by (c.f. Eq. 60)
@@
Figure 3: Predicted two-dimensional rephasing spectra from our theoretical model of an anyonic dimer system. (a) Real part of the third-order response in the normal regime ($\xi=0$); (b) Symmetry-broken regime ($\xi=1$), both computed at low temperature. (c, d) Evolution of the $\omega_{t}=\omega_{\tau}$ lineshape as the statistical angle $\theta$ varies from 00 to $\pi$. Dashed overlays track the real parts of the normal mode frequencies, illustrating bifurcation across exceptional points. These theoretical results highlight spectral signatures that may be accessible via coherent nonlinear spectroscopy.
In Fig. 2(a,b) we show the total relaxation rates for the anyon dimer as we vary the statistical angle from 0 to $\pi$ and the correlation parameter, $-1\leq\xi\leq 1$. Parameters are chosen to correspond to physically reasonable values of the hopping integral and relaxation rates. (Note for $\xi<0$, dashed and solid curves are swapped.) For $\theta<2$, both normal modes of the systems relax at the same rate. However, for $\theta>2$ the two rates bifurcate and concurrently the two frequencies coalesce, indicating the presence of an exceptional point that moves towards higher values of $\theta$ as the correlation parameter weakens. At high temperature, we see both an upper and lower critical value of $\theta$ for a given $\xi$. On the bosonic side ($\theta<\pi/2$), the relaxation rate is dominated by the purely thermodynamic contribution as per Eq. 8. As temperature increases, the lower point moves towards higher values of the statistical angle as $\xi$ increases. On the fermionic side, $\theta>\pi/2$, the statistical contribution dominates and the upper critical value shifts towards the left. At very high temperatures, the two meet at $\theta=\pi/2$ suggesting that Majorana quasi-particles are insensitive to environmental correlation effects.
@@ II.3 Detection via non-linear spectroscopy
Coherent multidimensional spectroscopy offers a powerful means to resolve the correlations and dynamics of anyonic systems, where exchange statistics deviate from standard bosonic or fermionic behavior. Unlike linear techniques, which conflate homogeneous and inhomogeneous broadening, coherent probes using phase-locked pulse sequences can isolate intrinsic dephasing and reveal pathways that may be sensitive to fractional statistics.[20, 21, 22] Focusing our attention on the rephasing third-order response, we write
@@
where $\tilde{\mathcal{G}}(\omega)$is the resolvent of the Liouvillian and $\mathcal{G}(t_{2})=e^{-i\mathcal{L}t_{2}}$ remains in the time domain and $\mu^{(\pm)}$ are superoperators representing the light-matter interaction. The $\pm$ indicates whether they act to the left or right and $\rho_{\text{eq}}$ is the equilibrium density matrix.[23] However, we have to point out that the dipole interaction $\mu^{(\pm)}$ also carries information concerning the statistical phase. This is because in a two-site system, where $a_{1}^{\dagger}$, $a_{2}^{\dagger}$ are creation operators for modes 1 and 2, the anyon dipole operator can be written as
@@ III Discussion
Fractional statistics deform the dissipative dynamics of coupled oscillators, allowing tunable mode protection and coherence control. Unlike bosonic or fermionic systems, anyonic oscillators have phase-sensitive decoherence rates and exceptional points that scale with temperature. This suggests a strategy to use exchange statistics and environmental correlations to create dissipation-resilient states in topological quantum devices. Our findings demonstrate that the exchange phase of anyons can be leveraged as a tunable parameter for modulating decoherence pathways, suggesting a new approach to coherence control via statistical engineering in open topological platforms. Further, we argue that coherent spectroscopic probes offer a wealth of detailed insight into the subtle interplay between the statistical phase and correlations within the environment.
The bifurcation between symmetric and antisymmetric mode lifetimes observed in our model reflects a dynamical symmetry-breaking process analogous to spontaneous $\mathcal{PT}$-symmetry breaking in non-Hermitian quantum systems. When the environmental coupling becomes strongly correlated ($\xi\to\pm 1$), the Liouvillian spectrum exhibits exceptional points—non-Hermitian degeneracies at which both eigenvalues and eigenvectors coalesce—leading to the emergence of dominant, long-lived modes. This phenomenon manifests as synchronization or antisynchronization of the coupled oscillators, depending on the sign of the bath correlation. The anyonic exchange phase enhances this symmetry breaking by modulating the off-diagonal terms in the effective dynamical matrix $W_{\mathrm{eff}}$, thereby shifting the location and character of the exceptional points. These findings connect naturally to broader efforts to characterize noise-induced phase locking in open quantum systems [16, 15, 24], and they echo the non-Hermitian transitions studied in optical and condensed matter platforms [25]. We believe that fractional statistics provide a novel and tunable axis for controlling dissipative phase transitions and coherence lifetimes in topological matter.
@@ Acknowledgements.
Supporting Information includes full derivations of the Lindblad dynamics, analytical forms of the statistical decay rate, and a Mathematica notebook reproducing all plots and simulations. Mathematica notebooks for the figures are available online at this Wolfram Cloud link. An accompanying WolframLanguage package is available at this link.
@@
Appendix A Dissipation rate for single oscillator
@@
$\frac{d\rho}{dt}=-i[H,\rho]+L_{\text{em}}\rho L_{\text{em}}^{\dagger}-\frac{1}{2}\{L_{\text{em}}^{\dagger}L_{\text{em}},\rho\}+L_{\text{abs}}\rho L_{\text{abs}}^{\dagger}-\frac{1}{2}\{L_{\text{abs}}^{\dagger}L_{\text{abs}},\rho\}.$ (17)
@@
$\mathcal{D}^{\dagger}_{\text{em}}[a]$ $=\frac{1}{2}\gamma(n_{\theta}+1)\{a,[a,a^{\dagger}]\},$ (22)
$\mathcal{D}^{\dagger}_{\text{abs}}[a]$ $=-\frac{1}{2}\gamma n_{\theta}\{a,[a,a^{\dagger}]\},$ (23)
@@
$\mathcal{D}^{\dagger}[a]=\frac{1}{2}\gamma\left[(n_{\theta}+1)-n_{\theta}\right]\{a,[a,a^{\dagger}]\}=\frac{\gamma}{2}\{a,[a,a^{\dagger}]\}.$ (24)
Assuming anyonic commutation relations $[a,a^{\dagger}]=\Phi(N)\approx\langle e^{iN\theta}\rangle$, the evolution becomes:

### v3

@@ I Introduction
We recently investigated noise-induced synchronization in open quantum systems consisting of two coupled quantum oscillators interacting with a common, correlated dissipative environment [14, 15, 16] where we assume that the environmental noise channels $\mathbf{E}_{1}(t)$ and $\mathbf{E}_{2}(t)$ evolve according to a coupled Ornstein–Uhlenbeck (OU) process written in differential form:
[…]
Here, the parameter $\xi\in[-1,1]$ controls the degree of instantaneous correlation between the two noise channels. The matrix $\mathbf{B}$ determines the noise amplitude and, together with $\boldsymbol{\Xi}$, sets the stationary variance of the process. In the limit of perfect environmental correlation ($\xi=\pm 1$), the oscillators exhibit long-lived phase synchronization, particularly when they are spectrally near-resonant. The presence of nonzero quantum discord between subsystems indicated that correlated environmental noise can generate persistent quantum coherences even in the absence of direct coupling. We further showed that two spins can synchronize their relative phases through shared stochastic driving, akin to the escapement mechanism that regulates motion in classical pendulum clocks.
@@ II Theoretical Development
We begin by developing a Lindblad framework for anyonic oscillators with fractional-statistics, where the algebra of creation and annihilation operators is deformed by an exchange phase $\theta$ that interpolates continuously between bosonic and fermionic limits. For a single anyon oscillator, the operators $a^{\dagger}$ and $a$ satisfy the $q$-deformed commutation relation:
@@ II.1 Anyon master equation
We first consider the dynamics of a single anyon oscillator coupled to a dissipative bath using the adjoint Lindblad formalism. The system is governed by the Hamiltonian $H=\hbar\omega a^{\dagger}a$, where $a$ and $a^{\dagger}$ are the anyon annihilation and creation operators, respectively. Dissipative coupling is introduced via two Lindblad jump operators: $L_{\text{em}}=\sqrt{\gamma(n_{\theta}+1)}\,a$ for spontaneous and stimulated emission, and $L_{\text{abs}}=\sqrt{\gamma n_{\theta}}\,a^{\dagger}$ for absorption. The thermal occupation number
$n_{\theta}=\frac{1}{e^{\beta\omega}-e^{i\theta}}$
may be complex due to the anyonic exchange phase $\theta$.
[…]
@@
which remains finite at $\theta=0$ (c.f. Fig. 1).
For bosonic systems ($\theta=0$), the statistical contribution vanishes, leaving only the thermal component. As the anyon phase increases towards the fermonic regime, the stastistical contribution increase monotonically until its maximum value at $\theta=\pi$ for a purely fermonic system. Consequently, moving from a purely fermionic system towards a fractionally quantized system results in a decrease in relaxation rate of the oscillator at a given temperature.
@@ II.2 Anyon double oscillator
with normal mode frequencies $\omega_{\pm}=\omega\pm J\cos(\theta)$. Following along the lines of our previous works[14, 15, 16] we define Lindblad jump operators in terms of the deformed normal mode operators as
@@
with explicit coefficients $\lambda_{k}^{(\pm)}$ for the emission and absorption channels are given in the Appendix B. Using these, we derive the equations of motion for the deformed anyonic operators which we write in matrix form as
@@
where the effective evolution matrix is given by
@@
Figure 3: Predicted two-dimensional rephasing spectra from our theoretical model of an anyonic dimer system. (a) Real part of the third-order response in the normal regime ($\xi=0$); (b) Symmetry-broken regime ($\xi=1$), both computed at low temperature. (c, d) Evolution of the $\omega_{t}=\omega_{\tau}$ lineshape as the statistical angle $\theta$ varies from 0 to $\pi$. Dashed overlays track the real parts of the normal mode frequencies, illustrating bifurcation across exceptional points. These theoretical results highlight spectral signatures that may be accessible via coherent nonlinear spectroscopy.
In Fig. 2(a,b) we show the total relaxation rates for the anyon dimer as we vary the statistical angle from 0 to $\pi$ and the correlation parameter, $1\leq\xi\leq 1$. Parameters are chosen to correspond to physically reasonable values of the hopping integral and relaxation rates. (Note for $\xi<0$, dashed and solid curves are swapped.) For $\theta<2$, both normal modes of the systems relax at the same rate. However, for $\theta>2$ the two rates bifurcate and concurrently the two frequencies coalesce, indicating the presence of an exceptional point that moves towards higher values of $\theta$ as the correlation parameter weakens. At high temperature, we see both an upper and lower critical value of $\theta$ for a given $\xi$. On the bosonic side ($\theta<\pi/2$), the relaxation rate is dominated by the purely thermodynamic contribution as per Eq. 10. As temperature increases, the lower point moves towards higher values of the statistical angle as $\xi$ increases. On the fermionic side, $\theta>\pi/2$, the statistical contribution dominates and the upper critical value shifts towards the left. At very high temperatures, the two meet at $\theta=\pi/2$ suggesting that Majorana quasi-particles are insensitive to environmental correlation effects.
@@ II.3 Detection via non-linear spectroscopy
Coherent multidimensional spectroscopy offers a powerful means to resolve the correlations and dynamics of anyonic systems, where exchange statistics deviate from standard bosonic or fermionic behavior. Unlike linear techniques, which conflate homogeneous and inhomogeneous broadening, coherent probes using phase-locked pulse sequences can isolate intrinsic dephasing and reveal pathways that may be sensitive to fractional statistics.[20, 21, 22]
In a typical three-pulse 2D spectroscopy experiment, the system is driven by a sequence of phase-stabilized laser pulses with well-defined wave vectors $\mathbf{k}_{1}$, $\mathbf{k}_{2}$, and $\mathbf{k}_{3}$. The emitted nonlinear signal appears in specific directions determined by the phase-matching condition, which enforces conservation of photon momentum among the optical fields. For the rephasing signal–often associated with a photon echo–the signal wave vector is given by $\mathbf{k}_{\text{sig}}=-\mathbf{k}_{1}+\mathbf{k}_{2}+\mathbf{k}_{3}$. This particular combination isolates Liouville pathways that involve evolution with negative time-ordering in the first interval, enabling partial cancellation of inhomogeneous broadening. It is important to emphasize that these wave vectors refer to the directions of the applied laser fields and are not related to any Bloch wave vectors or crystal momentum within the sample. The phase-matching geometry provides a versatile means of isolating specific dynamical pathways in the nonlinear response, independent of the underlying translational symmetry. For a comprehensive discussion of these techniques and phase-matching rules, we refer the reader to the foundational texts by Mukamel [23] and Hamm and Zanni [24].
Focusing our attention on the rephasing third-order response, we write
@@
where $\tilde{\mathcal{G}}(\omega)$ is the resolvent of the Liouvillian, with $\omega_{t}$ and $\omega_{\tau}$ being the Fourier-conjugate variables of $t_{3}$ and $t_{1}$ respectively, and $\mathcal{G}(t_{2})=e^{-i\mathcal{L}t_{2}}$ remains in the time domain and $\mu^{(\pm)}$ are superoperators representing the light-matter interaction. The $\pm$ indicates whether they act to the left or right and $\rho_{\text{eq}}$ is the equilibrium density matrix.[23] However, we have to point out that the dipole interaction $\mu^{(\pm)}$ also carries information concerning the statistical phase. This is because in a two-site system, where $a_{1}^{\dagger}$, $a_{2}^{\dagger}$ are creation operators for modes 1 and 2, the anyon dipole operator can be written as
@@ III Discussion
Fractional statistics deform the dissipative dynamics of coupled oscillators, allowing tunable mode protection and coherence control. Unlike bosonic or fermionic systems, anyonic oscillators have phase-sensitive decoherence rates and exceptional points that scale with temperature. This suggests a strategy to use exchange statistics and environmental correlations to create dissipation-resilient states in topological quantum devices.
Our findings demonstrate that the exchange phase of anyons can be leveraged as a tunable parameter for modulating decoherence pathways, suggesting a new approach to coherence control via statistical engineering in open topological platforms. Further, we argue that coherent spectroscopic probes offer a wealth of detailed insight into the subtle interplay between the statistical phase and correlations within the environment.
The bifurcation between symmetric and antisymmetric mode lifetimes observed in our model reflects a dynamical symmetry-breaking process analogous to spontaneous $\mathcal{PT}$-symmetry breaking in non-Hermitian quantum systems, as the eigenvalues of the dynamical matrix are no longer complex conjugate symmteric after symmetry breaking. When the environmental coupling becomes strongly correlated ($\xi\to\pm 1$), the Liouvillian spectrum exhibits exceptional points—non-Hermitian degeneracies at which both eigenvalues and eigenvectors coalesce—leading to the emergence of dominant, long-lived modes. This phenomenon manifests as synchronization or antisynchronization of the coupled oscillators, depending on the sign of the bath correlation. The anyonic exchange phase enhances this symmetry breaking by modulating the off-diagonal terms in the effective dynamical matrix $W_{\mathrm{eff}}$, thereby shifting the location and character of the exceptional points. These findings connect naturally to broader efforts to characterize noise-induced phase locking in open quantum systems [16, 15, 25], and they echo the non-Hermitian transitions studied in optical and condensed matter platforms [26]. We believe that fractional statistics provide a novel and tunable axis for controlling dissipative phase transitions and coherence lifetimes in topological matter.
@@ Acknowledgements.
Appendix includes full derivations of the Lindblad dynamics, analytical forms of the statistical decay rate, and a Mathematica notebook reproducing all plots and simulations. Mathematica notebooks for the figures are available online at this Wolfram Cloud link. An accompanying WolframLanguage package is available at this link.
@@
Appendix A Single oscillator dynamics
This section provides detailed derivations of key expressions appearing in the main text, including the statistical contribution to the relaxation rate, the adjoint form of the Lindblad dissipator for a single anyonic oscillator, and the functional form of the exchange-phase expectation value $\langle e^{i\theta N}\rangle$. These results underlie the coherence dynamics of the system and clarify how the anyonic exchange statistics modify both thermal occupation and dissipation. While the main text presents compact formulas suitable for presentation, we include these step-by-step derivations here to assist readers seeking deeper insight into the underlying calculations. In particular, we aim to support reviewers and readers who may be unfamiliar with statistical-phase averaging or with how Lindblad dynamics are extended to systems obeying deformed commutation relations.
A.1 Anyon Commutation Structure
We first recall the definition of the deformed algebra:
$aa^{\dagger}-e^{i\theta}a^{\dagger}a$ $=1$
So the commutator,
$[a,a^{\dagger}]$ $=aa^{\dagger}-a^{\dagger}a=1+e^{i\theta}N-N,$
[…]
where $\hat{N}=a^{\dagger}a.$ is the number operator. In the bosonic limit $(\theta=0)$:
$\Phi(\hat{N})=1.$
In the fermionic limit ($\theta=\pi$):
$\Phi(\hat{N})=1-2\hat{N}$.
A.2 Statistical Phase Averaging
To capture the effect of anyonic exchange statistics on coherence decay, we compute the thermal expectation of the phase factor
$\langle e^{i\theta\hat{N}}\rangle=\sum_{n=0}^{\infty}e^{i\theta n}\,P_{n},$
where $P_{n}=(1-z)z^{n}$ is the canonical probability for occupying the n-th level in a harmonic mode with $z=e^{-\beta\hbar\omega}$. Recognizing the sum as a geometric series, we obtain the closed-form expression
$\langle e^{i\theta\hat{N}}\rangle=\frac{1-z}{1-ze^{i\theta}},$
whose real part becomes
$\text{Re}\langle e^{i\theta\hat{N}}\rangle=\frac{(1-z)(1-z\cos\theta)}{1-2z\cos\theta+z^{2}}.$
As we show below, the statistical contribution to the relaxation rate then naturally follows as
[…]
highlighting how the thermal population distribution and anyonic phase combine to suppress coherence.
A.3 Heisenberg equations of motion
We begin with the diagonal Hamiltonian for a single anyonic mode,
$H=\hbar\omega\,\hat{N},\quad\text{where }\hat{N}=a^{\dagger}a.$
The Heisenberg equation of motion for the annihilation operator $a$ is given by
$\frac{da}{dt}=\frac{i}{\hbar}[H,a]=i\omega[\hat{N},a].$
To evaluate this, we use the identity
$\hat{N},a]=a^{\dagger}[a,a]+[a^{\dagger},a]a.$
[…]
$\hat{N},a]=-\Phi(\hat{N})a.$
Hence,
$\frac{d}{dt}a=-i\omega\Phi(\hat{N})a.$
[…]
However, for the purposes of deriving the relaxation rates, we approximate
$H,a]\approx-\hbar\omega a,$
corresponding to the bosonic case where $\Phi(\hat{N})=1$. This approximation allows us to isolate the effects of dissipation and fractional statistics in the Lindblad term without introducing nonlinearities into the coherent evolution. The validity of this approximation is justified in the regime where the deformation from bosonic statistics is modest (i.e., $\theta\ll\pi$) or when the statistical effects enter dominantly through the dissipator. Numerical simulations confirm that this substitution introduces negligible error in the computed coherence decay rates for the physical parameter ranges considered.
A.4 Dissipative terms
@@
$\frac{d\rho}{dt}=-i[H,\rho]+L_{\text{em}}\rho L_{\text{em}}^{\dagger}-\frac{1}{2}\{L_{\text{em}}^{\dagger}L_{\text{em}},\rho\}$
$+L_{\text{abs}}\rho L_{\text{abs}}^{\dagger}-\frac{1}{2}\{L_{\text{abs}}^{\dagger}L_{\text{abs}},\rho\}.$ (19)
@@
$\mathcal{D}^{\dagger}_{\text{em}}[a]$ $=\frac{1}{2}\gamma(n_{\theta}+1)\Phi(\hat{N})a,$ (24)
$\mathcal{D}^{\dagger}_{\text{abs}}[a]$ $=-\frac{1}{2}\gamma n_{\theta}\Phi(\hat{N})a,$ (25)
@@
$\mathcal{D}^{\dagger}[a]=\mathcal{D}_{\text{em}}^{\dagger}+\mathcal{D}_{\text{abs}}^{\dagger}=\frac{1}{2}\gamma\Phi(\hat{N})a$ (26)
[…]
@@
Here, we assume the unitary contribution remains linear in $a$ as discussed in Section 1.3, and we retain the full statistical correction via the dissipator.
@@
Remark. The adjoint form of the Lindblad dissipator, $\mathcal{D}^{\dagger}[A]$, is valid strictly under the trace operation, i.e., when computing expectation values of observables via $\langle A\rangle=\mathrm{Tr}[\rho\,A]$. It is derived to satisfy the dual relation $\mathrm{Tr}[\rho\,\mathcal{D}^{\dagger}[A]]=\mathrm{Tr}[(\mathcal{D}[\rho])\,A]$. Outside of the trace, $\mathcal{D}^{\dagger}[A]$ may not preserve the algebraic structure of $A$, especially in systems with deformed commutation relations. For this reason, care must be taken when interpreting $\mathcal{D}^{\dagger}$ as an operator-valued object rather than a generator of expectation value dynamics.
A.4.1 Limiting Cases: Bosons and Fermions
Bosonic limit ($\theta=0$). In the limit $\theta\to 0$, the anyonic algebra reduces to the canonical bosonic commutation relation $a,a^{\dagger}]=1$. The statistical average becomes:
$\langle e^{i\theta N}\rangle\to 1,\quad\text{so}\quad\Gamma_{\text{stat}}\to 0,$
and the full relaxation rate reduces to the thermal contribution alone:
$\Gamma_{\text{full}}^{\text{(boson)}}=\frac{\gamma}{2}(2n+1),$
where $n=\langle N\rangle=\frac{1}{e^{\beta\hbar\omega}-1}$. This result matches the textbook expression for the decay rate of $\langle a\rangle$ in a damped bosonic mode coupled to a thermal bath (see Breuer & Petruccione, Carmichael, Gardiner & Zoller).
Fermionic limit ($\theta=\pi$). In the fermionic limit, the algebra becomes $a,a^{\dagger}]=1-2\hat{N}$ and the occupation number is $n_{F}=\langle N\rangle=\frac{1}{e^{\beta\hbar\omega}+1}$. The statistical phase average is:
$\langle e^{i\pi N}\rangle=\sum_{n}(-1)^{n}P_{n}=1-2n_{F},$
so the statistical contribution becomes:
$\Gamma_{\text{stat}}=\frac{\gamma}{2}\left(1-(1-2n_{F})\right)=\gamma n_{F}.$
The thermal part becomes $\frac{\gamma}{2}(2n_{F}+1)$, yielding:
$\Gamma_{\text{full}}^{\text{(fermion)}}=\frac{\gamma}{2}(2n_{F}+1)+\gamma n_{F}=\frac{\gamma}{2}(2n_{F}+1)+\gamma n_{F}.$
This simplification reflects the exclusion principle, where $n_{F}\leq 1$ and higher occupation is forbidden. The resulting decay rate is reduced compared to the bosonic case at low temperatures but saturates at high $T$.
Summary. These limiting cases validate the general anyonic expression for $\Gamma_{\text{full}}$ and illustrate how the statistical contribution smoothly interpolates between enhanced decoherence (bosons) and suppressed coherence decay (fermions) depending on $\theta$ and $T$.